40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 12 00:59:45 2011, ryan, rana, Update, PSL, Return of the PSL temperature box 
    Reply  Sat Jul 30 19:06:25 2011, ryan, rana, Update, PSL, Return of the PSL temperature box rctempbox.png
       Reply  Sat Jul 30 20:41:50 2011, rana, Update, PSL, RefCav Stabilization back on Untitled.png
Message ID: 4697     Entry time: Thu May 12 00:59:45 2011     Reply to this: 5071
Author: ryan, rana 
Type: Update 
Category: PSL 
Subject: Return of the PSL temperature box 

The PSL temperature box has returned to service, with some circuit modifications. The 1k resistors on all the temp. sensor inputs (R3, R4, R7, R8, R12, R12) were changed to 0 Ohm. Also, the 10k resistors R26, R28, R29, and R30 were changed to 10.2k metal film. The DCC document will be updated shortly. There is now an offset in the MINCOMEAS channel compared to the others, which will be corrected in the morning after looking at the overnight trend.

ELOG V3.1.3-